The Official Report for Week Ending February 19, 2012

This forum is dedicated to the preservation of the WWII submarine USS Torsk in Baltimore, MD

Moderators: Gilbert, Bruce

The Official Report for Week Ending February 19, 2012

Postby Gilbert » Fri Mar 02, 2012 8:03 am

Summary of shipboard & offsite work for the week ending February 19, 2012

Saturday, February 18

Torsk Crew on board: Chief Mike, Dave, Frank, Ski, Butch, Cathy, Tracey, Michelle, Tim, John, Little Al, Nat, Gil, Charlie, Tristan and Jaime
Dave, Butch, Ski, Earl, Gilbert, Mike, and later Paul inspected the Conning Tower hatch to determine what needed to be done to repair it.

Chief Mike went to work troubleshooting the 1MC problems. First problem: the general alarm test switch did not work. Tested continuity through the pushbutton and got intermittent continuity. He took alcohol, put it in a spray bottle and applied liberally to each of the control switches and pushbuttons. While spraying, he cycled each of the switches to clean the contacts. This did the trick. The test circuit for each of the alarms (general, collision, and dive) all work. Next he tested each of the tubes in the signal generator. Found one bad tube, but unfortunately, we don’t have any spares in stock. This tube is the rectifier that supplies the DC bias and plate voltage to the signal generator amplifiers, so it’s no wonder the signal generator wasn’t producing any output.

John worked with Ski trouble shooting the topside SP phones. He spent the remainder of the day removing and reinstalling equipment that blocked the SP phone switch board. This was one of those jobs that involved work 2 minutes and then stand aside for 5 while visitors walked through, repeat, etc. Cathy helped with the project.

Dave helped load the three shelving units he picked up from Lowes during the week into LLAB for Gilbert and Bruce to organize various archival things. Frank, Tracey and Dave repaired the visitor damage to the Wardroom door hinges and secured the space. Dave worked with Butch on the measurements and design for installing an AC motor on the FER main engine LO pump. Butch will manufacture an adapter plate to allow use of the original pump mount.

Ski did a survey of topside SP Phone devices. After John and Ski tried to put a signal trace on the posts, they came up short on any signal coming into the boat. Still in the prelim stage. Ski took home two magneto's one (original ships) and another (strip ships version) to see if the later versions of this growler part can be configured to mount in the same position inside the original boxes and still have the crank arm line up.

Michelle stripped some paint in ATR.

Frank removed ganged rheostat from port cubicle to fabricate a new connection to the control wheel

Tim polished and buffed in engine rooms, started restoration of panels and relief valves in AER starboard forward

Little Al took photos and replaced light bulb.

Tracy did some general housekeeping that included changing a bulb in LLER, polishing valve handles throughout the boat and one propeller on a torpedo in ATR.

Gil worked in Archives.

Charlie, Tristan and Nat did housekeeping.

Offsite during the Week:

Gil worked on TVA Admin duties

*****************
Onsite Volunteer Hours – 106
Offsite Volunteer Hours - 3
Total Week Hours – 109

2012 Year to Date Volunteer Hours – 621

That’s it for the week.
Gil Bohannon
Gilbert
 
Posts: 601
Joined: Thu Nov 30, 2006 8:15 pm
Location: Not too far from Torsk

Return to General Discussion

Who is online

Users browsing this forum: No registered users and 11 guests

cron